OpenStudy (anonymous):

Helppp? Which of the following is the slope between the two points (-3, 1) and (3, -6)?

OpenStudy (nurali):

\[slope=\frac{ y2-y1 }{ x2-x1 }\]

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

Yes i already tried that, but it wont come out as one of the options given!

hartnn (hartnn):

1-(-6) =1+6=... = numerator -3-3 =...=denominator

OpenStudy (anonymous):

the options are 6/7 -6/7 7/6 -7/6

hartnn (hartnn):

1-(-6) =1+6=7= numerator -3-3 =-6=denominator so, -7/6

OpenStudy (anonymous):

Oh okayy that makes sense to me now. thanks!!

hartnn (hartnn):

welcome ^_^